OpenStudy (anonymous):

PLEASE HELP ME WITH A FEW QUESTION!!!! please show your work. 1)12^0t^7u^-11 simplify 2)8(3*10^14) Simplify write in scientific notation 3)(4c^4)(ac^3)(-3a^5c) Simplify 4)(9*10^7) (3*10^16) Simplify scientific notation 5)p(p^-7q^3)^-2q^-3 Simplify 6)3^2(3x)^3 simplify

OpenStudy (anonymous):

7)10m^6n^3 over 5M^2n^7 simplify 8)an investment of $5000 doubles in value every decade. The function f(x) = 5000 * 2^x, where x is the number of decade, models the growth of the value of the investment. How much is the investment worth after 30 year? Evaluate function and give value. 9)y = 25,600 * 1.01^x identify the initial amount of A and the growth factor of b in the exponential function 10)y = 2* 0.68 does this represent exponential growth, decay, or neither?

OpenStudy (anonymous):

do you have any ideas

OpenStudy (anonymous):

I do have some ideas I just need some help working through the problems.

OpenStudy (anonymous):

I think youll have a better chance getting answers if you post one question at a time

OpenStudy (anonymous):

can you help me on any of them?

OpenStudy (anonymous):

I will give medals to the best responses!